Write the equation of the line in slope-intercept form with slope of -3 passing through the point (6, -2).

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16946)

equation of line in slope point form
(y-y1)=m(x-x1) point (6,-2) slope -3
(y-6)=-3(x-(-3))
y-6 = -3x -9
y= -3x-3



Answer by ikleyn(52818)

.
Write the equation of the line in slope-intercept form with slope of -3 passing through the point (6, -2).
~~~~~~~~~~~~~~~~~


            An experimental fact is that @mananth is almost  ALWAYS  wrong.

            He (or she) is really almost  ALWAYS  wrong,  excluding that rare cases,  when he  (or she)  is right.

            This time,  he  (or she)  is wrong,  too.

            So I came to bring a correct solution.


equation of line in slope point form

(y-y1) = m(x-x1)  point (6,-2) slope -3

(y - (-2) ) = -3(x-6)

y + 2 = -3x + 18

y= -3x + 16          ANSWER

Solved.

----------------

It is very elementary and basic skill that every student must develop to the level to solve such problems on his  (or her)
own  AUTOMATICALLY  without asking assistance from outside.

    Similar to as you answer  "4"   when you are asked  "How many is  2*2  ?".

To help you to develop such skills,  special lesson was created in this site

    - Equation for a straight line having a given slope and passing through a given point

Find there a complete explanation on how to do it and learn the subject once and for all.


Consider this lesson as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.